the perimeter of a rectangle is 264 and the length is 72. find the width
Answer by stanbon(48535)

the perimeter of a rectangle is 264 and the length is 72. find the width
-------------------------------------------------------------
Perimeter = 2(L+W)
264 = 2(72+W)
132 = 72+W
W = 60
==============
Cheers,
Stan H.